# Information geometry of Lévy processes and financial models

### v2

@@ 3. Information geometry of Lévy processes
(75) $\mathcal{J}\propto(\det{g_{ij}})^{1/2}=\bigg(\det{\Big(-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i},\tilde{\partial}_{j})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}}\Big)}\bigg)^{1/2}.$
This Jeffreys prior can also be employed as a baseline for constructing Bayesian predictive priors. According to Komaki [25], the following Bayesian predictive priors provide improved predictive performance over the Jeffreys prior:
@@ 4.1. Tempered stable processes
We present the geometric structures of tempered stable processes, which form a subclass of Lévy processes. Choi [8] found the information geometry of various tempered stable processes and its statistical applications in parameter estimation and Bayesian prediction. Since the examples used in Choi [8] include generalized tempered stable (GTS) process, classical tempered stable (CTS) processes and rapidly-decreasing temepered stable processes, we compare our results with those previously established in the literature.
A tempered stable distribution is characterized by a Lévy triplet of the form $(0,\nu,\gamma)$ [30], that the Lévy measure $\nu$ is represented with
@@
Eq. (83) is consistent with the $\alpha$-divergence presented in Choi [8]. Specifically, when $\alpha=\pm 1$, the divergence coincides with the Kullback–Leibler divergence and its dual divergence found in the literature. Moreover, for $\alpha\neq\pm 1$, the $\alpha$-divergence in Choi’s work corresponds to the linear term in $T$ from Eq. (51) which is also related to Eq. (59). These expressions closely agree when $T$ is sufficiently small or when the parameters of $\mathbb{P}$ and $\mathbb{Q}$ are close enough.
@@
It is noteworthy that all other parameters of two GTS processes, except for $\lambda_{\pm}$, are identical due to the existence condition of the Radon–Nikodym derivative in Proposition 1. More details on this condition can be found in Rachev et al. [30] and Choi [8].
@@
As mentioned above, the $\alpha$-divergences for $\alpha=\pm 1$ are matched with those in Choi [8]. The $\alpha$-divergence for $\alpha\neq\pm 1$ in Choi [8] is the linear term in $T$ of Eq. (95). This is also related to Eq. (59).
@@
where $i,j$ run for $\lambda_{+}$ and $\lambda_{-}$. Moreover, the non-vanishing components of $\alpha$-connection in GTS geometry are derived as
@@
The other components are vanishing. When $\alpha=0$, Eq. (99) and Eq. (100) correspond to Levi-Civita connection which is also given in Choi [8].
It is worth noting that the Fisher information matrix and the $\alpha$-connection for GTS processes coincide with those derived in Choi [8]. Moreover, the same study introduced various statistical applications of GTS geometry, including penalized likelihood estimation for bias reduction [17, 26, 27] and Bayesian predictive priors [25]. Since our geometric results are equivalent, the same conclusions apply to these statistical methodologies.
@@ 4.2. CTS processes and CGMY model
By applying the CTS condition from Eq. (103) to the GTS Lévy measure in Eq. (87), the Lévy measure of CTS processes is represented with
@@
As mentioned for GTS processes, the existence of the Radon–Nikodym derivative under Proposition 1 requires that all parameters of two CTS processes, except for $\lambda_{\pm}$, are identical. Additional details regarding this condition can be found in the literature [30, 8].
@@
It is also evident that the same result can be obtained directly from Theorem 1. Moreover, similar to the case of GTS processes, the $\alpha$-divergence in Eq. (110) coincides with the results in Choi [8] when $\alpha=\pm 1$. For $\alpha\neq\pm 1$, the expression in Choi [8] corresponds to the linear term in $T$ from Eq. (110), also related to Eq. (59).
@@
where $i,j$ run for $\lambda_{+}$ and $\lambda_{-}$. Similarly, the non-trivial components of the $\alpha$-connection for CTS processes are given as
@@
All other components vanish. It is worth noting that the geometry derived directly from Theorem 2 coincides with the results in Eq. (113), Eq. (114), and Eq. (115).
@@
Moreorever, we can find potential candidates for a superharmonic function $\rho$ from the geometry for Komaki’s Bayesian predictive priors [25]. One example is as follows:
@@
where $c_{+}>0$ and $c_{-}>0$.
@@ 4.3. Variance gamma processes
Variance gamma (VG) processes [28] are also widely used in financial modeling. The VG process is a Lévy process characterized by the triplet $(0,\nu,\gamma)$. However, it does not fall within the class of tempered stable processes.
@@
However, the integral of Eq. (120) over $(-\infty,\infty)$ diverges making it ill-defined for computing the $\alpha$-divergence directly. To regularize the expression, we introduce a small positive parameter $a$ and consider the modified Lévy measure:
(121) $\nu(dx;\boldsymbol{\xi})=C\Big(\frac{{\rm e}^{-\lambda_{+}x}}{x^{1+a}}\mathbbm{1}_{x>0}(x)+\frac{{\rm e}^{-\lambda_{-}|x|}}{|x|^{1+a}}\mathbbm{1}_{x<0}(x)\Big)dx,$
This modified measure is integrable over the real line, allowing the use of standard divergence calculations, and the original measure can be recovered by taking the limit $a\to 0$.
It is noteworthy that Eq. (121) coincides with the Lévy measure of the CTS distribution, Eq. (104). This structural similarity suggests that one could formally obtain the VG process as a limiting case of the CTS process by setting $a=0$. While this makes it tempting to directly substitute $a=0$ into the $\alpha$-divergence expression for CTS processes, Eq. (110), caution is necessary as the Gamma function is undefined at zero. Additionally, $a=0$ is not allowed for CTS and GTS processes.
@@
By applying Eq. (124) to Eq. (110) and taking the limit as $a\to 0$, we obtain the $\alpha$-divergence between two variance gamma processes as:
(129) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-TC\big(\big(-\frac{1-\alpha}{2}\log{\lambda_{+}}-\frac{1+\alpha}{2}\log{\tilde{\lambda}}+\log{(\frac{1-\alpha}{2}\lambda_{+}+\frac{1+\alpha}{2}\tilde{\lambda}_{+})\big)}\\ +\big(-\frac{1-\alpha}{2}\log\lambda_{-}-\frac{1+\alpha}{2}\log\tilde{\lambda}_{-}+\log(\frac{1-\alpha}{2}\lambda_{-}+\frac{1+\alpha}{2}\tilde{\lambda}_{-})\big)\big)\big]\Big\}&(\alpha\neq\pm 1)\\ TC\Big(\big(\tilde{\lambda}_{+}/\lambda_{+}-1+\log{\lambda_{+}}-\log{\tilde{\lambda}_{+}}\big)+\big(\tilde{\lambda}_{-}/\lambda_{-}-1+\log{\lambda_{-}}-\log{\tilde{\lambda}_{-}}\big)\Big)&(\alpha=-1)\\ TC\Big(\big(\lambda_{+}/\tilde{\lambda}_{+}-1+\log{\tilde{\lambda}_{+}}-\log{\lambda_{+}}\big)+\big(\lambda_{-}/\tilde{\lambda}_{-}-1+\log{\tilde{\lambda}_{-}}-\log{\lambda_{-}}\big)\Big)&(\alpha=1)\end{array}\right.$.
By taking the limit of $a\to 0$ to Eq. (113), the Fisher information matrix of variance gamma processes is represented with
@@
where $i,j,$ and $k$ run for $\lambda_{\pm}$. The $\alpha$-connection of variance gamma process geometry is given as
@@
All other components are vanishing.
As with GTS and CTS processes, the information geometry of variance gamma processes enables a range of statistical applications. Since the VG geometry is also e-flat, we can consider the penalized likelihood method for bias reduction [17]. Based on the Fisher information matrix for variance gamma processes given by Eq. (132), the corresponding Jeffreys prior is obtained as
@@
This Jeffreys prior can be leveraged to construct Bayesian predictive priors, as discussed in Komaki [25]. Moreover, superharmonic functions for variance gamma processes can be obtained by taking the limit $a\to 0$ in the corresponding superharmonic functions for CTS processes:
@@ 5. Conclusion
In this paper, we have developed an information-geometric framework for Lévy processes, significantly extending precedent works focused on the information geometry of tempered stable processes [8] and the Kullback–Leibler divergence of Lévy processes [14]. By deriving the $\alpha$-divergence for Lévy processes, we constructed the corresponding Fisher information matrices and $\alpha$-connections, uncovering the underlying differential-geometric structure of this broad and important class of stochastic processes. Statistical advantages from obtaining the information geometry were also discussed.
Moreover, we demonstrated the utility of this geometric framework by applying it to several prominent Lévy models widely used in finance, including tempered stable processes, CGMY models, and variance gamma processes. For each of these models, we characterized their information geometry and discussed potential applications to bias-reduced estimation and Bayesian predictive priors that are tools relevant to model calibration, estimation robustness, and risk measurement in financial contexts. These examples underscore how information geometry functions not only as a theoretical framework but also as a practical methodology in quantitative modeling.
In closing, this work contributes to the growing interdisciplinary study of stochastic process theory, information geometry, and applied probability. By embedding Lévy processes within the information-geometric framework, we offer a novel foundation for analyzing and comparing complex models exhibiting jumps and non-Gaussian behavior. We hope that this study inspires further research at the intersection of geometry, statistics, and stochastic modeling.

### v3

@@ 3. Information geometry of Lévy processes
(76) $\mathcal{J}\propto(\det{g_{ij}})^{1/2}=\bigg(\det{\Big(-\frac{4}{1-\alpha^{2}}\Delta^{(\alpha)}_{T}(\partial_{i}||\tilde{\partial}_{j})|_{\tilde{\boldsymbol{\xi}}=\boldsymbol{\xi}}\Big)}\bigg)^{1/2}.$
This Jeffreys prior can also be employed as a baseline for constructing Bayesian predictive priors. According to Komaki [25], the following Bayesian predictive prior $\tilde{\mathcal{J}}$ provides improved predictive performance over the Jeffreys prior:
@@ 4.1. Tempered stable processes
We start with the geometric structures of tempered stable processes, which form a subclass of Lévy processes. Choi [8] found the information geometry of various tempered stable processes and its statistical applications in parameter estimation and Bayesian prediction. Since the examples used in Choi [8] include generalized tempered stable (GTS) process, classical tempered stable (CTS) processes and rapidly-decreasing tempered stable processes, we compare our results with those previously established in the literature.
A tempered stable distribution is characterized by a Lévy triplet of $(0,\nu,\gamma)$ [31], that the Lévy measure $\nu$ is in the form of
@@
It is noteworthy that Eq. (84) is consistent with the $\alpha$-divergence presented in Choi [8]. Specifically, when $\alpha=\pm 1$, the divergence coincides with the Kullback–Leibler divergence and its dual divergence found in the literature. Moreover, for $\alpha\neq\pm 1$, the $\alpha$-divergence in the literature corresponds to the linear term in $T$ from Eq. (52) which is also related to Eq. (60). These expressions closely agree when $T$ is sufficiently small or when the parameters of $\mathbb{P}$ and $\mathbb{Q}$ are close enough.
@@
It is noteworthy that all the parameters of two GTS processes, except for $\lambda_{\pm}$, are identical due to the existence condition of the Radon–Nikodym derivative in Proposition 1. More details on this condition can be found in Rachev et al. [31], Kim and Lee [21], and Choi [8].
@@
As mentioned above, the $\alpha$-divergences for $\alpha=\pm 1$ are matched with those in Choi [8]. The $\alpha$-divergence for $\alpha\neq\pm 1$ in the literature is the linear term in $T$ of Eq. (96). This is also related to Eq. (60).
@@
where $i,j$ run for $\lambda_{+}$ and $\lambda_{-}$. Moreover, the $\alpha$-connection of GTS geometry is derived as
@@
The other components are vanishing. It is easy to check that the manifold is e-flat.
It is worth noting that the Fisher information matrix and the $\alpha$-connection for GTS processes coincide with those derived in Choi [8]. Moreover, the literature introduced various statistical applications of GTS geometry, including penalized likelihood estimation for bias reduction [17, 26, 27] and Bayesian predictive priors [25]. Since the geometric results found here are equivalent, the same conclusions apply to these statistical methodologies.
@@ 4.2. CTS processes and CGMY model
By applying the CTS condition of Eq. (104) to the GTS Lévy measure in Eq. (88), the Lévy measure of CTS processes is represented with
@@
As mentioned for GTS processes, the existence of the Radon–Nikodym derivative under Proposition 1 requires that all the parameters of two CTS processes, except for $\lambda_{\pm}$, are identical. Additional details regarding this condition can be found in the literature [31, 21, 8].
@@
It is also evident that the same result can be obtained directly from Theorem 1. Moreover, similar to the case of GTS processes, the $\alpha$-divergence in Eq. (111) coincides with the results in Choi [8] when $\alpha=\pm 1$. For $\alpha\neq\pm 1$, the expression in the literature corresponds to the linear term in $T$ from Eq. (111), also related to the linear version of the $\alpha$-divergence given by Eq. (60).
@@
where $i,j$ run for $\lambda_{+}$ and $\lambda_{-}$. Similarly, the $\alpha$-connection of CTS process geometry is given by
@@
All other components vanish. It is easy to check that the geometry is e-flat. It is obvious that the geometry derived directly from Theorem 2 coincides with the results in Eq. (114), Eq. (115), and Eq. (116).
@@
This Jeffreys prior can be used for the penalized log-likelihood in Eq. (75).
Moreover, we can find Komaki’s Bayesian predictive priors [25] by searching potential candidates for a superharmonic function $\rho$ from the geometry. One example for a superharmonic function is as follows:
@@
where $c_{+}>0$ and $c_{-}>0$. These superharmonic functions are plugged to Eq. (77) for Bayesian predictive priors.
@@ 4.3. Variance gamma processes
Variance gamma (VG) processes [28] are also widely used in financial modeling. The VG process is a Lévy process characterized by a triplet $(0,\nu,\gamma)$. However, it does not fall within the class of tempered stable processes.
@@
However, the integral of Eq. (121) over $(-\infty,\infty)$ diverges, making it ill-defined for computing the $\alpha$-divergence directly. To regularize the expression, we introduce a small positive parameter $a$ and consider the modified Lévy measure:
(122) $\nu(dx;\boldsymbol{\xi})=C\Big(\frac{{\rm e}^{-\lambda_{+}x}}{x^{1+a}}\mathbbm{1}_{x>0}(x)+\frac{{\rm e}^{-\lambda_{-}|x|}}{|x|^{1+a}}\mathbbm{1}_{x<0}(x)\Big)dx.$
This modified measure is integrable over the real line, allowing the use of standard divergence calculations, and the original measure can be recovered by taking the limit of $a\to 0$.
It is noteworthy that Eq. (122) coincides with the Lévy measure of the CTS distribution, Eq. (105). This structural similarity may suggest that one could formally obtain the VG process as a limiting case of the CTS process by setting $a=0$. While this makes it tempting to directly substitute $a=0$ into the $\alpha$-divergence expression for CTS processes, Eq. (111), extra caution is necessary as the Gamma function is undefined at zero. Additionally, $a=0$ is not allowed for CTS and GTS processes.
@@
By applying Eq. (125) to Eq. (111) and taking the limit of $a\to 0$, we obtain the $\alpha$-divergence between two variance gamma processes as
(130) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big\{1-\exp\big[-TC\big(\big(-\frac{1-\alpha}{2}\log{\lambda_{+}}-\frac{1+\alpha}{2}\log{\tilde{\lambda}}+\log{(\frac{1-\alpha}{2}\lambda_{+}+\frac{1+\alpha}{2}\tilde{\lambda}_{+})\big)}\\ +\big(-\frac{1-\alpha}{2}\log\lambda_{-}-\frac{1+\alpha}{2}\log\tilde{\lambda}_{-}+\log(\frac{1-\alpha}{2}\lambda_{-}+\frac{1+\alpha}{2}\tilde{\lambda}_{-})\big)\big)\big]\Big\}&(\alpha\neq\pm 1)\\ TC\big((\tilde{\lambda}_{+}/\lambda_{+}-1+\log{\lambda_{+}}-\log{\tilde{\lambda}_{+}})+(\tilde{\lambda}_{-}/\lambda_{-}-1+\log{\lambda_{-}}-\log{\tilde{\lambda}_{-}})\big)&(\alpha=-1)\\ TC\big((\lambda_{+}/\tilde{\lambda}_{+}-1+\log{\tilde{\lambda}_{+}}-\log{\lambda_{+}})+(\lambda_{-}/\tilde{\lambda}_{-}-1+\log{\tilde{\lambda}_{-}}-\log{\lambda_{-}})\big)&(\alpha=1)\end{array}\right.$.
From the $\alpha$-divergence of Eq. (130), the Fisher information matrix of variance gamma processes is represented with
@@
where $i,j,$ and $k$ run for $\lambda_{\pm}$. The same metric tensor is also derived from taking the limit of $a\to 0$ to Eq. (114).
Similarly, the $\alpha$-connection of variance gamma process geometry is given by
@@
All other components are vanishing. It is easy to check that the variance gamma geometry is e-flat.
Similar to the GTS and CTS processes, the information geometry of variance gamma processes enables a range of statistical applications. Since the VG geometry is also e-flat, we can consider the penalized likelihood method for bias reduction [17]. Based on the Fisher information matrix for variance gamma processes given by Eq. (133), the corresponding Jeffreys prior is obtained as
@@
This Jeffreys prior for variance gamma processes can be leveraged to construct Bayesian predictive priors, as discussed in Komaki [25]. Moreover, superharmonic functions for variance gamma processes can be obtained by taking the limit of $a\to 0$ in the corresponding superharmonic functions for CTS processes:
@@
4.4. Merton model
The Merton model [29] extends the Black-Scholes framework by incorporating discontinuous jumps in asset prices through a combination of geometric Brownian motion and a Poisson jump component. The model simulates not only routine market fluctuations but also abrupt price changes. The consideration on those components provides a more robust explanation for the volatility smile and the mispricing of out-of-the-money options than the standard Black-Scholes model.
Unlike the tempered stable processes and variance gamma processes studied earlier, the Merton model is a Lévy process with a Lévy triplet of $(\sigma,\nu,\gamma)$, i.e., $\sigma\neq 0$. Its Lévy measure is given by
(140) $\nu(dx)=\frac{\lambda}{\delta\sqrt{2\pi}}\exp{\Big(-\frac{(x-m)^{2}}{2\delta^{2}}\Big)}dx,$
where $\lambda,\delta>0$ and $m\in\mathbb{R}$.
Let us consider two risk-neutral Merton models with the same volatility of $\sigma$. Since the Merton model is a risk-neutral exponential Lévy model, the $\alpha$-divergence between two Merton models can be derived from Eq. (48) in Theorem 1:
(146) $D^{(\alpha)}(\mathbb{P}||\mathbb{Q})=\left\{\begin{array}[]{ll}\frac{4}{1-\alpha^{2}}\Big(1-{\rm{e}}^{-\Delta^{(\alpha)}_{T}(\mathbb{P}||\mathbb{Q})}\Big)&(\alpha\neq\pm 1)\\ \frac{T}{2\sigma^{2}}\big(\lambda(\exp{(m+\frac{\delta^{2}}{2})}-1)-\tilde{\lambda}(\exp{(\tilde{m}+\frac{\tilde{\delta}^{2}}{2})}-1)\big)^{2}+&\\ T\big((\tilde{\lambda}-\lambda)+\lambda(\ln{(\frac{\lambda}{\delta})}-\ln{(\frac{\tilde{\lambda}}{\tilde{\delta}})}-\frac{1}{2}+\frac{(m-\tilde{m})^{2}+\delta^{2}}{2\tilde{\delta}^{2}})\big)&(\alpha=-1)\\ \frac{T}{2\sigma^{2}}\big(\lambda(\exp{(m+\frac{\delta^{2}}{2})}-1)-\tilde{\lambda}(\exp{(\tilde{m}+\frac{\tilde{\delta}^{2}}{2})}-1)\big)^{2}+&\\ T\big((\lambda-\tilde{\lambda})+\tilde{\lambda}(\ln{(\frac{\tilde{\lambda}}{\tilde{\delta}})}-\ln{(\frac{\lambda}{\delta})}-\frac{1}{2}+\frac{(m-\tilde{m})^{2}+\tilde{\delta}^{2}}{2\delta^{2}})\big)&(\alpha=1)\end{array}\right$.,
where $\Delta_{T}^{(\alpha)}(\mathbb{P}||\mathbb{Q})$ is given by
$\Delta_{T}^{(\alpha)}(\mathbb{P}||\mathbb{Q})$ $=\frac{1-\alpha^{2}}{4}\frac{T}{2\sigma^{2}}\big(\lambda(\exp{(m+\frac{\delta^{2}}{2})}-1)-\tilde{\lambda}(\exp{(\tilde{m}+\frac{\tilde{\delta}^{2}}{2})}-1)\big)^{2}$
$\quad+T\Big(\frac{1-\alpha}{2}\lambda+\frac{1+\alpha}{2}\tilde{\lambda}-\Big(\frac{\lambda}{\delta}\Big)^{\frac{1-\alpha}{2}}\Big(\frac{\tilde{\lambda}}{\tilde{\delta}}\Big)^{\frac{1+\alpha}{2}}(\frac{1-\alpha}{2}\frac{1}{\delta^{2}}+\frac{1+\alpha}{2}\frac{1}{\tilde{\delta}^{2}})^{-1/2}$
$\exp{\big[\frac{1}{2}(\frac{1-\alpha}{2}\frac{1}{\delta^{2}}+\frac{1+\alpha}{2}\frac{1}{\tilde{\delta}^{2}})^{-1}(\frac{1-\alpha}{2}\frac{m}{\delta^{2}}+\frac{1+\alpha}{2}\frac{\tilde{m}}{\tilde{\delta}^{2}})^{2}-\frac{1}{2}(\frac{1-\alpha}{2}\frac{m^{2}}{\delta^{2}}+\frac{1+\alpha}{2}\frac{\tilde{m}^{2}}{\tilde{\delta}^{2}})\big]}\Big).$
It is straightforward to check that the Kullback–Leibler divergence in Eq. (146) is identical to that in Cont and Tankov [14].
From Eq. (63) in Theorem 2 or Eq. (69) in Corollary 2, the Fisher information matrix of the Merton model is given by
(147) $g_{ij}=T\begin{pmatrix}\frac{1}{\lambda}+\frac{(e^{m+\delta^{2}/2}-1)^{2}}{\sigma^{2}}&\frac{\lambda(e^{m+\delta^{2}/2}-1)e^{m+\delta^{2}/2}}{\sigma^{2}}&\frac{\lambda\delta(e^{m+\delta^{2}/2}-1)e^{m+\delta^{2}/2}}{\sigma^{2}}\\ \frac{\lambda(e^{m+\delta^{2}/2}-1)e^{m+\delta^{2}/2}}{\sigma^{2}}&\frac{\lambda}{\delta^{2}}+\frac{\lambda^{2}e^{2(m+\delta^{2}/2)}}{\sigma^{2}}&\frac{\lambda^{2}\delta e^{2(m+\delta^{2}/2)}}{\sigma^{2}}\\ \frac{\lambda\delta(e^{m+\delta^{2}/2}-1)e^{m+\delta^{2}/2}}{\sigma^{2}}&\frac{\lambda^{2}\delta e^{2(m+\delta^{2}/2)}}{\sigma^{2}}&\frac{2\lambda}{\delta^{2}}+\frac{\lambda^{2}\delta^{2}e^{2(m+\delta^{2}/2)}}{\sigma^{2}}\end{pmatrix}$,
where the coordinate system is $(\lambda,m,\delta)$.
Similarly, the $\alpha$-connection for information geometry of the Merton model is derived from Eq. (64) in Theorem 2 or Eq. (70) in Corollary 2:
[…]
Similar to the other processes in this paper, statistical benefits of information geometry can be discussed. First of all, the penalized likelihood estimation is not available for the Merton model because none of the Merton model geometry nor its submanifolds are e-flat. However, we still can search the Bayesian predictive priors of the Merton model although the metric tensor and the Jeffreys priors of the Merton model are much more complicated than the other processes we covered earlier.
@@ 5. Conclusion
In this paper, we have developed an information-geometric framework for Lévy processes, significantly extending precedent works focused on the information geometry of tempered stable processes and the Kullback–Leibler divergence of Lévy processes. By deriving the $\alpha$-divergence of Lévy processes, we constructed the Fisher information matrix and the $\alpha$-connection for this broad and important class of stochastic processes. Statistical advantages from obtaining the information geometry were also discussed.
Moreover, we demonstrated this geometric framework of several prominent Lévy models, widely used in finance, including tempered stable processes, the CGMY model, variance gamma processes, and the Merton model. For these models, we characterized their information geometry and discussed potential applications to bias-reduced estimation and Bayesian predictive priors that are tools relevant to model calibration, estimation robustness, and risk measurement in the contexts of quantitative finance. These examples exhibit how information geometry functions not only as a theoretical framework but also as a practical methodology in quantitative modeling.
In conclusion, this work contributes to the growing interdisciplinary study of stochastic process theory, information geometry, applied probability, and quantitative finance. By embedding Lévy processes within the information-geometric framework, we offer a novel foundation for analyzing and comparing complex models with jumps and non-Gaussian behavior. We hope that this study inspires further research at the intersection of geometry, statistics, stochastic processes and financial modeling.